OpenStudy (anonymous):

How do you do this the format is y-y1=m(X-X1) The problem: the line passes through points at (-2,-2) and (-1,-6) Whats the answer?

OpenStudy (anonymous):

first find the slope. m = y-y/x-x

OpenStudy (anonymous):

it is eight negative three. what would that reduce to?

OpenStudy (anonymous):

m= (-6)-(-2)/ (-1)-(-2) = -6+2/ -1+2 = -4/1 =-4

OpenStudy (anonymous):

I got -4 as the slope.

OpenStudy (anonymous):

so then you choose one of the points and plug it in for y1 and x1

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

either works because they are both points on the line. (-2,-2) will be easier mathwise

OpenStudy (anonymous):

thank you. you and gingersnapz helped me alt..